A force f̂ is expressed with respect to the basis ê1 and ê2 by equation 2(3 ê1 +4 ê2) N. If the directions of ê1 makes an angle of 30 degree with f̂, find the vector ê2.


1
Expert's answer
2021-01-26T13:16:14-0500

Given,

"f=6\hat{e_1}+8 \hat{e_2}"

The direction "\hat{e_1}" makes angle with "=30^\circ"

"\Rightarrow \tan(30^\circ)=|\frac{8\hat{e_2}}{6\hat{e_2}}|"

"\Rightarrow |\frac{1}{\sqrt{3}}|=|\frac{8\hat{e_2}}{6\hat{e_1}}|"

squaring both side,

"\Rightarrow \frac{1}{3}=\frac{64\hat{e_2^2}}{36\hat{e}^2}"

"\Rightarrow 36 \hat{e_1^2}=3\times64\hat{e_2^2}"

"\Rightarrow \hat{e_1^2}=\frac{16}{3}\hat{e_2^2}"